### v3

# Barycenter kinematics in Local Group analogues

Abstract
We present a new way to identify systems similar to our Local Group (LG) of galaxies in cosmological simulations. Our method uses as a new constraint the speed and direction of the LG’s center of mass, which we can measure accurately from cosmic microwave background data. When we apply these criteria to different cosmological simulations and compare the results with traditional selection methods, we find statistically significant differences. Our approach produces simulated galaxy pairs where the relative M31 velocity is less radial ($2\%$ to $7\%$ difference over the mean) and more tangential ($1\%$ to $3\%$ difference over the mean) than in cases that do not take into account the barycenter speed. The radial change pattern appears consistently across all cosmological models we test, showing that matching the observed barycenter velocity has a measurable effect when modeling and interpreting Local Group-like systems.
keywords:
Local Group,N-body simulations, barycenter kinematics
1 Introduction
Understanding the orbit between M31 and the Milky Way (MW) is key to explaining how the Local Group (LG) evolved (Hartl & Strigari, 2024). Recent research challenges the old view that this orbit is mostly head-on (radial), suggesting instead that it has more sideways motion than previously thought (van der Marel et al., 2019; Salomon et al., 2021). These studies rely on cosmological simulations to establish expectations for the tangential velocity components. To study LG dynamics, previous work typically identify similar galaxy pairs in simulations by applying specific criteria to halo pairs, including their masses, approach speed, distance, and isolation from other large structures (Forero-Romero et al., 2013; Sawala et al., 2023; Hartl & Strigari, 2022).
We identified an important measurement that previous studies haven’t used when defining LG-like systems. This measurement is the LG’s barycenter velocity ($\mathbf{v}_{b}$), which we can determine precisely from cosmic microwave background (CMB) data (Planck Collaboration et al., 2020). We use this constraint in two ways. First, we consider its magnitude ($v_{b}$). Second, we examine the direction, measured as the angle ($\theta$) between the barycenter velocity vector and the position vector of M31 relative to the MW ($\mathbf{r}$), as shown in Figure 1. We express this angle more conveniently as $\mu=\cos\theta$, which comes from the dot product of the unit vectors for $\mathbf{v}_{b}$ and $\mathbf{r}$.
Figure 1: Diagram showing the alignment between key quantities used to select LG analogues. $\mathbf{v}_{b}$ represents the barycenter velocity of the LG and $\mathbf{r}$ shows the position of M31 relative to the MW.
Previous studies couldn’t use the barycenter velocity because they worked with simulations that were too small in volume. Recently, Forero-Romero & Sierra-Porta (2022) showed that peculiar velocities can be inaccurate in small simulation boxes due to power spectrum limitations. Using simulations with volumes ranging from 100 to 2967 Mpc across, they demonstrated that accurate kinematic properties require boxes at least 1 Gpc across. Both theoretical arguments and simulation results support this conclusion.
In this paper, we use simulations to measure how including the barycenter velocity constraint affects the predicted radial and tangential velocities of M31 relative to the MW. We also explore how these effects vary across simulations with different cosmological parameters.
Our paper is organized as follows: Section 2 covers the observational constraints, Section 3 describes the simulations used and the different cosmological models. In Section 4, we explain how we select LG analogues using both traditional criteria and our new approach. Section 5 presents and discusses our findings, and Section 6 summarizes our conclusions.
Table 1: Summary of observed reference values. The columns represent, respectively: the physical observable, its measured value, the corresponding units, and the bibliographic source.
{ruledtabular}
Observable Value Units Source
$v_{b}$ $620\pm 15$ $\text{km\ s}^{-1}$ Planck Collaboration et al. (2020)
$(l,b)_{v_{b}}$ ($271.9\pm 2.0$, $29.6\pm 1.4$) deg Planck Collaboration et al. (2020)
$r_{\text{MW}}$ $8.29\pm 0.16$ kpc McMillan (2011)
$(l,b)_{\text{MW}}$ $(359.944215\pm 0.000006,-0.046079\pm 0.000009)$ deg. Yusef-Zadeh et al. (1999)
$r_{M31}$ $761\pm 11$ kpc Li et al. (2021)
$(l,b)_{\text{M31}}$ $(121.174409\pm 0.000022,-21.572996\pm 0.000022)$ deg. Evans et al. (2010)
$v_{r,\text{M31}}$ $109.3\pm 4.4$ $\text{kms}^{-1}$ van der Marel et al. (2012)
$v_{t,\text{M31}}$ $82.4\pm 31.2$ $\text{kms}^{-1}$ Salomon et al. (2021)
$\mu$ $0.88\pm 0.01$ adim. This work
2 Observational constraints
Table 1 summarizes the key observational values used to calculate $\mu$. We obtain the barycenter velocity of the LG from cosmic microwave background (CMB) measurements (Planck Collaboration et al., 2020). The positions of both M31 and the MW are well known from previous studies. For the Sun-to-Galactic Center distance, we use the value from McMillan (2011), who derived this by fitting MW mass models to observational data. The Galactic Center position comes from VLA measurements reported by Yusef-Zadeh et al. (1999). For M31, we use the distance measured by Li et al. (2021) using Cepheid variable stars, and coordinates from the Chandra Source Catalog (Evans et al., 2010).
To calculate the relative position vector from M31 to MW ($\mathbf{r}$) and determine $\mu$, we use a Monte Carlo approach. We build a covariance matrix using coordinate measurement error ellipses (from NASA/IPAC Extragalactic Database) and sample from a multivariate normal distribution. The uncertainties in Table 1 represent the standard deviations of these samples. Our method yields an M31-MW separation of $765\pm 11$ kpc, matching observations reported by Chamberlain et al. (2022). To compute $\mu$, we generate normal distribution samples for $\mathbf{v}_{b}$ and the distances to MW and M31. We then compute the dot product between $\mathbf{v}_{b}$ and $\mathbf{r}$, finding $\mu=-0.88\pm 0.01$. This result shows a strong anti-alignment between these vectors.
3 Simulation
For this study, we use the AbacusSummit simulation suite (Maksimova et al., 2021; Garrison et al., 2021). We analyze the base simulation box with $6912^{3}$ particles in a 2 Gpc/$h$ volume. Because the simulation snapshot is at redshift $z=0.1$, we use the halos’ peculiar velocities to project their positions to present-day values.
[…]
We also study ten reference cosmologies (c010, c012-c018, c021-022) that match models from other major projects, mostly using massless neutrinos. For wider parameter testing, we include simulations with a linear derivative grid (c100-c116) with small changes across eight parameters, and an unstructured emulator grid (c130-c181) covering broader parameter ranges. All cosmologies use the same $\tau$ value (0.0544) and most included 60 meV neutrinos, consistent with Planck 2018 results.
Figure 2: Left panel: cumulative distribution of $\mu$, the cosine of the angle between the barycenter velocity of the LG and the position vector of M31 relative to the MW. We also plot a uniform distribution from -1 to 1 for comparison Right panel: Cumulative distribution for the barycenter speed $v_{b}$ of the Local Group.
4 Selection of LG analogues
We use halos identified with the CompaSO Halo Finder (Hadzhiyska et al., 2021), with selection criteria similar to those in Forero-Romero & Sierra-Porta (2022). First, we selected halos with maximum circular velocities $V_{\text{max}}>200$ km s-1. From these, we identify isolated pairs as two halos, A and B, that are each other’s nearest neighbors. We also require that no third halo with a maximum circular velocity greater than halo B’s can be found within three times the pair’s separation distance. We define halo A as the one with the greater maximum circular velocity, which we designated as M31. This forms our main sample, typically containing about $8\times 10^{5}$ pairs per simulation box.
From this main sample, we create two sub samples:
The first subsample, which we call traditional, includes pairs meeting these conditions:
  • Total pair mass between $0.5,5.0]\times 10^{12}$ M
  • Separation between halos less than 1 Mpc $h^{-1}$
  • Negative radial relative velocity of M31 with respect to the MW ($v_{r}<0$ km s-1), after accounting for cosmic expansion
The second subsample, which we call realistic, includes all the traditional criteria plus these additional barycenter velocity conditions:
  • Barycenter speed greater than the observed value, including the uncertainty ($v_{b}>605$ km s-1)
  • Alignment value less than the observed value, including the uncertainty ($\mu<-0.87$)
We use inequalities rather than narrow ranges around the observed values because this approach: (1) captures the observed trends that the barycenter speed is unusually high and the alignment shows strong anti-alignment, and (2) allows us to keep more pairs in our sample that satisfy these conditions.
Nonetheless, we include for completeness a variation of the realistic subsample, hereafter referred to as centered, which replaces the one-sided bounds with a symmetric interval around the observed mean values:
  • 500 $^{-1}<v_{b}<740 kms-$1
  • $0.80<\mu<-0.96$
The traditional subsample typically contains about $2\times 10^{5}$ pairs per simulation box, while the realistic and centered sample contain about $10^{4}$ pairs.
4.1 Statistical Tests
Using the simulations, we compute the cumulative probability distributions for $\mu$ and three key Local Group properties:
  • Total Local Group mass ($M_{\text{LG}}$)
  • Radial velocity of M31 relative to the MW ($v_{r}$)
  • Tangential velocity of M31 relative to the MW ($v_{t}$)
To measure how our barycenter velocity constraints affect these properties, we compare the mean values between the traditional and realistic pair distributions. We also use two-sample Kolmogorov–Smirnov (KS) tests to determine if these differences are statistically significant, with $p$-values below 0.05 considered significant.
Figure 3: Mean velocity components of M31 relative to the MW as a function of $v_{b}$ and $\mu$ in the base cosmology (c000) for the traditional sample. The top panel shows $v_{r}$ and the bottom panel shows $v_{t}$
5 Results
5.1 Base cosmology
Figure 2 shows the cumulative distributions of $\mu$ and $v_{b}$ in the base cosmology (c000). These distributions illustrate the key features of our barycenter velocity constraints used to select LG analogues. The $\mu$ distribution of reveals that $\mathbf{v}_{b}$ and $\mathbf{r}$ in simulated traditional LG pairs tend to be either aligned or anti-aligned. We confirm this by comparing against a uniform distribution using a KS test, which shows statistically significant differences ($p<0.05$). Using the same statistical method, we also find that the distribution is not symmetric with a slight tendency towards alignment.
We find statistically significant differences between traditional and realistic pairs in M31’s radial velocity distribution. For radial velocity ($v_{r}$), the mean value increases by 5%, indicating a smaller radial speed. The differences in M31’s tangential velocity distribution and total mass ($M_{\text{LG}}$) are not statistically significant.
Figure 3 shows the mean $v_{r}$ as a function of $v_{b}$ and $\mu$ for the traditional pairs in the base cosmology. The figure indicates that smaller radial velocities correspond to higher values of $v_{b}$ and more extreme values of $\mu$ i.e. more aligned or anti-aligned pairs. In contrast, no clear relationship is observed for the tangential velocity.
5.2 All cosmologies
Figure 4: Mean values of M31 velocity distributions in all cosmologies (top panel shows $v_{r}$ and the bottom panel $v_{t}$). The main y-axis corresponds to the mean values of the velocities for traditional and realistic pairs. The secondary y-axis shows the percentage difference in mean values of realistic pair distributions compared to the mean of traditional pair distributions in each cosmology. Dashed lines and shaded regions represent the average and standard deviation of the mean values of traditional pair distributions for all cosmologies
Figure 5: Statistical significance of the differences between realistic and traditional selection criteria for radial velocity $v_{r}$ (top panel) and tangential velocity $v_{t}$ (bottom panel). The $p$-values shown are from two-sample Kolmogorov-Smirnov tests across all cosmologies.
Table 2: Number of cosmologies where the difference between traditional and realistic or centered pairs was statistically significant for each combination of observables.
{ruledtabular}
Observable N. cosmo. (Realistic) N. cosmo. (Centered)
None 6 47
$v_{r}$ only 65 34
$v_{t}$ only 1 9
$v_{r}$ and $v_{t}$ 12 3
$M_{\text{LG}}$ only 0 0
$v_{r}$ and $M_{\text{LG}}$ 7 0
$v_{t}$ and $M_{\text{LG}}$ 0 0
All three 2 0
Total 93 93
Our barycenter velocity constraints consistently affect the distributions of both $v_{r}$ and $v_{t}$ compared to traditional constraints across different cosmological models.
Figure 4 shows how the mean values of tangential and radial velocities change across all cosmologies. Realistic pairs consistently show less negative $v_{r}$, with mean differences between -2% to -7%. On the other hand, $v_{t}$ tends to increase for realistic pairs. The effect in this case is weaker, with changes between -1% and 3%.
This stronger effect in $v_{r}$ is also reflected in the significance tests. Figure 5 shows the p-values of the KS tests between traditional and realistic pairs velocity distributions. Among the 93 cosmologies we study, the differences in radial velocity distribution are significant in 86 models, while tangential velocity distribution show significant differences in 15 cosmologies. The effect on total mass ($M_{\text{LG}}$) is significant in only 9 cosmologies. Table 2 summarizes for each cosmology, which combinations of observables display statistically significant differences between traditional and realistic or centered pairs.
Table 3: Percentage differences in the distributions of the M31 velocity components ($v_{r}$ and $v_{t}$) and the total mass of the Local Group ($M_{\text{LG}}$). For each observable and sample type, the table reports the mean difference relative to the traditional sample, computed (i) using only the cosmologies where the difference is statistically significant and (ii) using all 93 cosmologies. The last column gives the number of cosmologies where the ks test indicates a significant difference between the distributions.
{ruledtabular}
Observable Sample Mean difference I (%) Mean difference II (%) Significant cosmologies (out of 93)
$v_{r}$ Realistic 4.64-4.64 4.51-4.51 86
Centered 2.83-2.83 2.14-2.14 37
$v_{t}$ Realistic $1.63$ $0.96$ 15
Centered 0.75-0.75 0.33-0.33 12
$M_{\text{LG}}$ Realistic 0.83-0.83 0.31-0.31 9
Centered - 0.06-0.06 0
Figure 6: Cumulative distributions of M31 velocities relative to the MW in cosmology 004 and cosmology 130, where we found the biggest difference in the mean between traditional and realistic pairs. In cosmology c004, the difference in both $v_{r}$ (-7%) and $v_{t}$ (3%) was statistically significant. In cosmology 130 only the difference in $v_{t}$ (3%) was statistically significant.
We calculated the average shift in the observables across cosmologies with significant differences, and to avoid selection bias, we repeated the procedure across all cosmologies. The results shown in table 3 further show the mentioned trend in the measured shift of the observables. The change in $v_{r}$ seen in the realistic sample is the largest and most robust across cosmologies. The mean change of roughly 4.6% in the significant subset remains almost the same (4.5%) when averaged over all cosmologies. The centered sample follows the same trend but with smaller magnitudes. Tangential velocity behaves differently. The mean difference in significant cosmologies is small (around 1.6% for the realistic sample), and when averaged over all cosmologies the value drops below 1%. For the centered sample the average change is negative (which correspond smaller velocities), but the magnitude is below 1% for the both averages (across significant cosmologies and across all cosmologies). $M_{\text{LG}}$ has the weakest effect, with changes below 1 percent and significant differences in only 9 cosmologies for the realistic sample and none for the centered.
[…]
Although the barycenter-based selection produces mean shifts that are statistically significant given the large sample sizes, the absolute shifts ($\sim$ a few km $s^{-1}$) are small compared to the typical velocity dispersion of LG analogues (tens of km $s^{-1}$); thus the effect is a modest kinematic refinement rather than a major physical change.
Figure 7: Relative importance of various cosmological parameters for predicting whether the difference in M31’s velocities between centered and traditional pairs is statistically significant. The left panel shows the parameter importance for $v_{r}$, and the right panel for $v_{t}$. Importance is calculated using a Random Forest algorithm.
Figure 8: Scatter plots of the mean difference in M31’s velocities between centered and traditional pairs versus the most important cosmological parameters for predicting whether a cosmology leads to significant differences. The left panel corresponds to $v_{r}$, and the right panel for $v_{t}$. For each parameter, $\rho$ is the calculated Pearson correlation coefficient with respect to the mean difference in the velocity.
5.3 Cosmological parameters
To explore which cosmological parameters most strongly influence whether the differences in $v_{r}$ and $v_{t}$ between traditional and realistic pairs are statistically significant, we trained a Random Forest classifier using the KS-test outcome (significant vs. not significant) as the target. However, the realistic filter introduces a class imbalance. For $v_{r}$, in 86 out of 93 cosmologies we found significant differences and for $v_{t}$, only in 11. Under these conditions, the classifier is dominated by the majority class and cannot reliably learn the minority class. This is reflected on the performance metrics obtained: 0.48 macro F1 score and weighted F1 score of 0.89.
For this reason we also considered the centered filter which defines a broader range around the mean observed values of $v_{b}$ and $\mu$. This filter produces a more balanced distribution of significant and non-significant cosmologies. In this case the F1 macro score was 0.55 and the weighted F1 score was 0.58. For this sample, the mean importance obtained for each parameter with a stratified K-fold validation is shown in figure 7. The most important parameter for determining significance for both velocity components was $A_{s}$ (scalar perturbation amplitude).
We then computed Pearson correlation coefficients between the mean velocity differences and individual cosmological parameters. Under the realistic filter, $v_{r}$ shows no measurable correlation, while $v_{t}$ shows a moderate negative correlation ($\rho\approx-0.45$) with both $A_{s}$ and $\sigma_{8}$. Under the centered filter however, the mean difference in $v_{r}$ exhibits moderate positive correlations ($\rho\approx 0.3$), with both $A_{s}$ and $\sigma_{8}$, while the correlation with $v_{t}$ remains unchanged.
The appearance of $A_{s}$ and $\sigma_{8}$ among the most influential parameters in the Random Forest Analysis, and their moderate correlations under the centered filter, suggests that the amplitude of matter fluctuations may play a role in setting how strongly barycenter kinematics affect Local Group analogue velocities. A higher fluctuation amplitude deepens gravitational potentials and may influence how coherent the barycenter motion is relative to halo orbits. Exploring these mechanisms, however, requires a dedicated physical study and lies beyond the scope of this work. The results presented here should therefore be regarded as exploratory, illustrating statistical trends rather than establishing physical causation.
6 Conclusion
In this work, we explored the barycenter velocity ($\mathbf{v}_{b}$) as a new kinematic constraint for defining Local Group (LG) analogues. This quantity has very low observational uncertainty but was not included in previous studies due to the need for sufficiently large simulation volumes to achieve numerical convergence in peculiar velocities (Forero-Romero & Sierra-Porta, 2022). Using simulations from the AbacusSummit project, with a box size of 2 Gpc $h^{-1}$, we have shown that including this constraint introduces a small but systematic shift in the predicted velocity distributions of M31. We demonstrated this effect by comparing LG analogues selected using our proposed constraints (realistic pairs) with those selected using traditional constraints (separation distance less than 1 Mpc $h^{-1}$ and negative radial velocity). Our results show that in the realistic pairs, M31 tends to have a smaller radial velocity $v_{r}$ and larger tangential velocity $v_{t}$ compared to the traditional pairs. We quantified this effect using the difference in means and assessed its significance through two-sample Kolmogorov-Smirnov tests.
This effect is consistent across cosmological models and is particularly robust for $v_{r}$. In our analysis of 93 different cosmologies, we found significant differences in the radial velocity distribution for 86 models, while the tangential velocity distribution showed significant differences in 15 cosmologies. The mean differences between realistic and traditional pairs were -4.5% for $v_{r}$ (less radial) and 1% for $v_{t}$ (more tangential), averaged across all cosmologies.
We also found that $A_{s}$ and $\sigma_{8}$ are the cosmological parameter that most strongly influences the significance of differences in radial and tangential velocity distributions. However, further work is needed to understand the influence of cosmological parameters on LG kinematics. Similarly, while we have demonstrated statistical differences in velocity distributions, a deeper exploration of the physical mechanisms explaining why barycenter kinematics correlate with orbital properties remains for future work. Assessing if these results change our understanding of the Local Group properties and history –for instance in light of studies such as Chamberlain et al. (2022), Hartl & Strigari (2024),Benisty & Mota (2025)– is also left for future work.
In summary, our results show that the kinematic constraints proposed in this work consistently influence the radial and tangential velocities of M31, with effects that are robust across cosmological models. Although small in magnitude, these systematic effects highlight that the LG’s bulk motion encodes useful information about its internal dynamics. Incorporating these new kinematic constraints should therefore be considered when defining LG analogues in future studies.
Acknowledgments
This work made use of the NASA/IPAC Extragalactic Database (NED), funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.